Father and son are together 88 years old. How old is each of them, when his father was 38 year old at the birth of his son.
Answer by stanbon(75887)

Father and son are together 88 years old. How old is each of them, when his father was 38 year old at the birth of his son.
-----
Equations:
f + s = 88 yrs
f = s + 38
-----
Substitute for "f" and solve for "s":
s+38 + s = 88
2s = 50
s = 25 yrs (the son's age now)
-----
Solve for "f":
f = 25+38
f = 63 yrs (the father's age now)
====================
Cheers,
Stan H.
====================